The weather news says there is a 30% chance of rain this weekend. Make an area to represent all possibilities. Find the probability that it will rain.
a. No days
b. One day
c. Two days

Answer by stanbon(48546)

The weather news says there is a 30% chance of rain this weekend. Make an area to represent all possibilities. Find the probability that it will rain.
The weekend is two days.
P(rain)=0.3
P(no rain)=0.7
----------------
a. P(No days)=0.7^2 = 0.49
b. P(One day)=2C1(0.3)(0.7) = 2(0.21)=0.42
c. P(Two days)=0.3^2=0.09
==================
Cheers,
Stan H.